If I have an unknown quantity of gas at a pressure of 1.2 atm, a volume of 31 liters, and a temperature of 87°C, how many moles of the gas do I have? *
1
Expert's answer
2020-04-18T07:00:16-0400

Answer on Question #109907, Chemistry, General Chemistry

If I have an unknown quantity of gas at a pressure of 1.2 atm, a volume of 31 liters, and a temperature of 87°C, how many moles of the gas do I have?


Solution:

According to Mendeleev-Clapeyron's equation:

pV = nRT

where n is number of moles; p - pressure; V is a volume; T - temperature which equals (273+87) = 360 K; R - universal gas constant.

Thus the n number we can define as:

n = pV/(RT)

Using given values:

n = 1.2*31/(0.082*360) = 1.26 (moles)


Answer: n = 1.26 moles.